Over what grade should we find the 5 best students of 37 in a test with an average of 81 and standard deviation of 7?
Answer by ewatrrr(24785)

 
Hi,
m= 81, sd = 7, n = 37, top 5 0r top 13.5 percentile
Normsinv(.135) = z = 1.102 = (x - 81)/7
1.102 = (x - 81)/7
7.7 + 81 = 88.7 Over an 89 test score